Smith and Sons Inc. Are determining the viability of a new product line. The new Product will Require a $280,000 piece of equipment. Shipping and installation will cost $20,000. The equipment has a 3-year tax life, and the allowed depreciation for such property are 33%, 45%, 15%, and 7% for years 1 through 4. Inventory will increase by $15,000, account payable increasing by $8,000 and account receivables increasing by $10,000. The Product line is expected to generate annual revenue of $106,000 per year, with cost of goods sold being 36,000 per year and other costs (Excluding Depreciation) of $12,000 per year. The Tax rate is 30 Percent, Annual interest expense is $11,000 per year, and the required return for this project is 12 Percent.

find the year 0 cash flow, FCF0

Find the year 2 cash flow, FCF2

Please show all work
Answer #1

rate positively .. let me know if you need any clarification..

Computation of depreciation =
Total cost = 280,000+20,000 300000
year 1 2 3
Depreciation rate 33.00% 45.00% 15.00%
Depreciation amt               99,000            135,000              45,000
change in working capital
inventory 15000
Account payable -8000
Account receivable 10000
Net working capital 17000
Computation of NPV
year 0 1 2
a initial investment           (300,000)
b Working capital              (17,000)
A Initial investment           (317,000)
operating cash flow
i Revenue            106,000            106,000
ii Cost =36000+12000              48,000              48,000
iii depreciation              99,000            135,000
iv=i-ii-iii Profit before tax            (41,000)            (77,000)
v=iv*30% Tax@ 30%            (12,300)            (23,100)
vi=iv-v Profit after tax            (28,700)            (53,900)
B=vi+iii operating cash flow              70,300              81,100
Year 0 Cash flow =           (317,000)
Year 2 cash flow =               81,100